OpenStudy (anonymous):

Find all solutions of cos x + sin x tan x = 2 in the interval [0,2pi)

hartnn (hartnn):

how about writing tan x as sin x/cos x ? did you try this? or anything ?

OpenStudy (anonymous):

\[\cos x + \sin x \tan x = 2 \rightarrow \cos x + \sin x \frac{ \sin x }{ \cos x } =2\]

OpenStudy (anonymous):

You can solve it by multipliying sin x by sin x and adding fractions

OpenStudy (anonymous):

Let me know what you get pls